# Quantum Illumination and Quantum Radar: A Brief Overview

### v2

@@ III Quantum Radar Types and Criteria
Regardless of the type of protocol utilised there a number of criteria that must be met in order for a practical quantum radar to operate with quantum advantage. Furthermore, since quantum radar is a continually evolving and emerging technology, it can be difficult to be sure whether a given result would qualify as a proof of principle demonstration. The criteria below set out the fundamental requirements for a quantum target detection protocol to constitute a quantum radar.
@@ IV.1.1 The radar equation
by which current state of the art classical radar systems and their performances are modelled.
@@ IV.1.2 The signal-to-noise ratio (SNR)
The (effective) collecting area of the receiver, $a_{\mathrm{r}}$, can be further expressed in terms of the receiver gain, $G_{\mathrm{r}}=4\pi a_{\mathrm{r}}/\lambda^{2}$, allowing us to rewrite Eq. (5) as
$\mathrm{SNR}=\frac{P_{\mathrm{tx}}G_{\mathrm{tx}}G_{\mathrm{r}}\lambda^{2}\sigma F^{4}}{(4\pi)^{3}R^{4}k_{B}TB_{\mathrm{n}}F_{\mathrm{n}}}.$ (6)
@@ IV.1.3 Hypothesis testing
  • Probability of mis-detection (Type II error), $P_{\mathrm{miss}}$,
@@ V.2 Idler storage
Suppose one’s means of idler storage has an associated efficiency $\eta_{\mathrm{I}}$ quantifying the proportion of idlers successfully stored for later recombination with the returning signal. Then, the QCB for QI (Eq. (15) would be modified as
@@ V.3.3 Sum-frequency generating receiver
The underlying idea is motivated by the fact that one can construct a unitary $\hat{U}$ which makes one the POVMs for optimal binary discrimination of two equally likely states, $\hat{U}\rho_{0}\hat{U}^{\dagger}$ and $\hat{U}\rho_{1}\hat{U}^{\dagger}$, simply a projection onto the vacuum. In light of this, SFG ultimately takes, as input, low-brightness signal-idler mode pairs with either zero or non-zero cross correlations, and outputs a single mode which is either a vacuum or non-vacuum, respectively. After multiple SFG cycles and FF, this signal is subjected to photon counting measurements capable of achieving the minimum error probability defined by the Helstrom bound.
This particular receiver design was later used in the determination of QI’s receiver operating characteristic (ROC) which gives the detection probability as a function of the false alarm probability [61]. Nevertheless, despite its theoretical capability of saturating the QI QCB, such a receiver remains physically out of reach due to limitations on current technology, requiring both single-photon nonlinearities as well as efficient idler storage.
@@ V.5 Issue with loss tolerance
QI is robust to noise but not to loss. It provides advantage in the task of distinguishing an entangled signal photon within a bright thermal background, but as long as the overall round-trip loss is limited [68]. While this loss may be limited in optical implementations with collimated Gaussian beams, it fails to be the case in the microwave regime where the typical wave is much more isotropic and therefore subject to high geometric loss. As a result, even in the forward path, the outgoing beam is subject to increased losses which are easily alleviated at other wavelengths.
@@ VII Further modifications of the QI protocol
There exists wide interest across the quantum technology community in the study of non-Gaussianity and, with respect to QI, how it may be introduced to enhance the existing protocol. While Gaussian quantum states, operations and detection schemes offer a relatively straightforward means of physical implementation and mathematical description, their use together often places limitations within the realms of potential quantum enhancements. Alternative approaches of making use of non-Gaussianity include, for example, probe states where non-Gaussianity is introduced via photon addition or photon subtraction [73, 74]. It has also been shown that non-Gaussian receiver designs, in the form of a non-linear amplifier (NLA), has the potential of enhancing the effective SNR of QI to the extend of extending the range of possible regimes for which a quantum advantage over coherent states can be obtained [75].

### v3

# thanks: Corresponding author

@@ III Quantum Radar Types and Criteria
Regardless of the type of protocol utilised, several criteria must be met for a practical quantum radar to operate with quantum advantage. Furthermore, since quantum radar is a continually evolving and emerging technology, it can be difficult to be sure whether a given result would qualify as a proof of principle demonstration. The criteria below set out the fundamental requirements for a quantum target detection protocol to constitute a quantum radar.
@@ IV.1.1 The radar equation
which is the fundamental model for classical radar systems. Note that the use of a single value for radar cross section is a simplification. This cross section varies with the frequency of the radar wave, the target’s orientation, and its specific configuration. These factors affect how radar waves interact with the target, leading to significant variability in radar return signals.
@@ IV.1.2 The signal-to-noise ratio (SNR)
The (effective) collecting area of the receiver, $a_{\mathrm{r}}$, can be further expressed in terms of the receiver gain, $G_{\mathrm{r}}=4\pi a_{\mathrm{r}}/\lambda^{2}$, with $\lambda$ being the wavelength. This allows us to rewrite Eq. (5) as
$\mathrm{SNR}=\frac{P_{\mathrm{tx}}G_{\mathrm{tx}}G_{\mathrm{r}}\lambda^{2}\sigma F^{4}}{(4\pi)^{3}R_{\mathrm{T}}^{4}k_{B}TB_{\mathrm{n}}F_{\mathrm{n}}}.$ (6)
@@ IV.1.3 Hypothesis testing
  • Probability of mis-detection (Type II error), $P_{\mathrm{md}}$,
@@ V.2 Idler storage
Suppose one’s means of idler storage has an associated efficiency $\eta_{\mathrm{I}}$ quantifying the proportion of idlers successfully stored for later recombination with the returning signal. Then, the QCB for QI [Eq. (15)] would be modified as
@@ V.3.3 Sum-frequency generating receiver
The underlying idea is motivated by the fact that one can construct a unitary $\hat{U}$ which makes one of the POVMs for optimal binary discrimination of two equally likely states, $\hat{U}\rho_{0}\hat{U}^{\dagger}$ and $\hat{U}\rho_{1}\hat{U}^{\dagger}$, simply a projection onto the vacuum. In light of this, SFG ultimately takes, as input, low-brightness signal-idler mode pairs with either zero or non-zero cross correlations, and outputs a single mode which is either a vacuum or non-vacuum, respectively. After multiple SFG cycles and FF, this signal is subjected to photon counting measurements capable of achieving the minimum error probability defined by the Helstrom bound.
This receiver design was later used in the determination of QI’s receiver operating characteristic (ROC) which gives the detection probability as a function of the false alarm probability [61], in the Neyman-Pearson setting of QI [62].
Nevertheless, despite its theoretical capability of saturating the QI QCB, such a receiver remains physically out of reach due to limitations on current technology, requiring both single-photon nonlinearities as well as efficient idler storage.
@@ V.5 Issue with loss tolerance
QI is robust to noise but not to loss. It provides advantage in the task of distinguishing an entangled signal photon within a bright thermal background, but as long as the overall round-trip loss is limited [69]. While this loss may be limited in optical implementations with collimated Gaussian beams, it fails to be the case in the microwave regime where the typical wave is much more isotropic and therefore subject to high geometric loss. As a result, even in the forward path, the outgoing beam is subject to increased losses which are easily alleviated at shorter wavelengths.
@@ VII Further modifications of the QI protocol
There exists wide interest across the quantum technology community in the study of non-Gaussianity and, with respect to QI, how it may be introduced to enhance the existing protocol. While Gaussian quantum states, operations and detection schemes offer a relatively straightforward means of physical implementation and mathematical description, their use together often places limitations within the realms of potential quantum enhancements. Alternative approaches of making use of non-Gaussianity include, for example, probe states where non-Gaussianity is introduced via photon addition or photon subtraction [74, 75]. It has also been shown that non-Gaussian receiver designs, in the form of a non-linear amplifier (NLA), has the potential of enhancing the effective SNR of QI to the extent of extending the range of possible regimes for which a quantum advantage over coherent states can be obtained [76].